Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-5x – 7y = 25.5
-2x + 10y = -9

Answer by stanbon(75887)

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-5x – 7y = 25.5
-2x + 10y = -9
-----------
Multiply thru 1st equation by -2
Multiply thru 2nd equation by -5
-------
10x + 14y = -51
10x - 50y = 45
----
Subtract and solve for "y":
64y = -96
y = -3/2
==============
Solve for "x":
10x + 14y = -51
10x + 14(-3/2) = -51
10x - 21 = -51
10x = 30
x = 3
---------
Cheers,
Stan H.
==========================